Jack usually mows his lawn in 5 hours. Mary can mow the same yard in 3 hours. how much time would it take for them to mow the lawn together?
Found 3 solutions by edjones, richwmiller, solver91311:
Answer by edjones(8007)

Jack does the 1/5 the job in 1 hr & Mary 1/3 in 1 hr.
1/5 + 1/3
= 3/15 + 5/15 = 8/15 together in 1 hr or 15/8 hrs it takes to do the job together.
.
Ed

Answer by richwmiller(17219)

In one hour Jack mows 1/5 of the lawn
In one hour Mary mows 1/3 of the lawn
so we would think that together they can mow in less than 3 hours
x/5+x/3=1
15/8 hrs.=1.875 hrs. =1:52.5 hrs. almost two hours


15/(8*5)+15/(8*3)=1
Jack mows 3/8 and Mary mows 5/8
3/8+5/8=1

Answer by solver91311(24713)



If A can do a job in x time periods, then A can do of the job in 1 time period. Likewise, if B can do the same job in y time periods, then B can do of the job in 1 time period.

So, working together, they can do



of the job in 1 time period.

Therefore, they can do the whole job in:



time periods.


John